Assume that in a perfectly competitive market the supply is P = 8 + Q and the demand is P = 48 - 3Q.

Find the consumer surplus variation caused by a per unit tax of 2.
Answer #1

When the imposition of a per unit tax is not specifically mentioned between buyers and sellers , it is considered to be imposed on sellers .

Supply function: P = 8+Q Demand function : P = 48 - 39 - - - - - - * - at equilibrium, Demand & Supply 48 - 3Q = 8+ P => 4Q =Now, due to a per unit tax of a the supply equation becomes P = 8+ (Q+2) = Q + 10 The supply curve shifts upward at S NOW, O